OpenStudy (anonymous):

A college student has set aside $240 for the rest of the school year to use the coin - operated laundry facility in his dormitory. Each time he uses the machines, it costs $7.50. Choose the equation that represents the amount remaining in his fund, f , after he has done laundry x times. Find the amount remaining in the fund after 12 trips to the laundry facility.

OpenStudy (anonymous):

f(x) = 240-7.5*x f(12) = 240 - 7.5*12 =150

OpenStudy (anonymous):

Thank you so much :D